Suppose that the amount of time one spends in a bank is exponentially distributed with mean 10 minutes. What is the probability that a customer will spend more than 15 minutes in the bank? What is the probability that a customer will spend more than 15 minutes in the bank given that he is still in the bank after 10 minutes?


1
Expert's answer
2021-08-31T13:18:49-0400

Let "X=" the amount of time a customer spends in a bank: "X\sim Exp(\lambda)."


"f(x)=\lambda e^{-\lambda x}, x>0"

"\lambda=\dfrac{1}{\mu}=\dfrac{1}{10}=0.1"


"P(X>15)=\displaystyle\int_{15}^{\infin}f(x)dx"

"=\lim\limits_{A\to\infin}\displaystyle\int_{15}^{A}0.1 e^{-0.1 x}dx"

"=\lim\limits_{A\to\infin}[-e^{0.1x}]\begin{matrix}
 A \\
 15
\end{matrix}=-0+e^{-0.1(15)}"

"=e^{-1.5}\approx0.22313"

An exponential distribution has a memory-less property, i.e the future probabilities are not affected by any past data.


"P(X>15|X>10)=P(X>10+5|X>10)"

"=P(X>5)=\lim\limits_{A\to\infin}\displaystyle\int_{5}^{A}0.1 e^{-0.1 x}dx"

"=\lim\limits_{A\to\infin}[-e^{0.1x}]\begin{matrix}
 A \\
 5
\end{matrix}=-0+e^{-0.1(5)}"

"=e^{-0.5}\approx0.60653"